The gravitational field strength at Earth's surface is about 9.81 N/kg [down]. What is the gravitational field strength exactly 1.6 Earth radii from Earth's centre?
1
Expert's answer
2020-04-17T10:19:40-0400
"g(r)=g(R)\frac{R^2}{r^2}"

"g(1.6R)=g(R)\frac{1}{1.6^2}=\frac{9.81}{1.6^2}=3.83\frac{m}{s^2}"